The question
-14=-20y-7x
10y+4=2x
My attempt
-20y-7=-14
10y-2x=-4 (×2)
-20y-7x=-4
20y-4x=-8
-11x=-12
This i know is wing but help me figure out how to do this correctly please

Answer by Alan3354(69443)

The question
-14=-20y-7x
10y+4=2x
---------- That's not a question. It's 2 equations.
My attempt
-20y-7=-14 ****** What happened to the x? (That's a question.)
10y-2x=-4 (×2)
-20y-7x=-4 *********** Why did -14 become -4 ?
20y-4x=-8
-11x=-12
This i know is wing but help me figure out how to do this correctly please